7. What is the mass of oxygen combine with 15.76 g of gold in a sample of gold oxide, AuO2?



1
Expert's answer
2021-10-11T06:15:52-0400

Molar mass of "O_2=16g\/mol"

Molar mass of "Au=196.96g\/mol"

Total molecular mass "=196.97+(16\times 2)=228.67g/mol"

Mass of Gold in the sample"=15.76g"

Mole ratio "=1:2"

"=15.76\times2=31.52"

"=" "228.67\over 31.52""=7.25g"